OpenStudy (anonymous):

Indefinite integral of (tan^5 x)*(sec^4 x)dx Can't seem to figure out which identities to substitute in...

OpenStudy (p0sitr0n):

\[\int\limits_{?}^{?}tanx^{5 }*secx^4\] something like this i suppose?

OpenStudy (amistre64):

undo a chain rule .....

OpenStudy (anonymous):

\[\int\limits_{}^{}\tan^{5} (x)\sec^{4}(x)dx\]

OpenStudy (p0sitr0n):

take tanx^2 +1 =secx^2

OpenStudy (anonymous):

oooooh

OpenStudy (p0sitr0n):

it becomes tanx^5 * (tanx^2+1)secx^2 (tanx^7+tanx^5)secx^2 you get two integrals tanx^7 * secx^2 + tanx^5 * secx^2

OpenStudy (anonymous):

And then substitue u=tan(x), du=sec^2(x), right? and it all simplifies to [tan^6(x)]/6 + [tan^8(x)]/8 + C. Thanks.